# latexOverwriting file

### v4

@@ 4 Discussion and Future Directions
We have provided the first analysis of the error incurred by ensemble Kalman filters, as approximations of true filtering distribution, beyond the linear Gaussian setting. We have employed an appropriate weighted TV metric and obtained new stability estimates in this metric, in order to establish the approximation results. Our framing of the problem is motivated by the framing of the analysis of the particle filter contained in [50, Section 1], a systematization of the original proof of convergence of the particle filter appearing in [14]. Although it introduces new methodology and theoretical results for nonlinear Kalman filters, our work leaves open numerous avenues for further analysis; we now highlight those that we identify as particularly important. These remaining open problems are substantial, but the framework we map out in this paper is an appropriate one in which to address them.
@@
    Our theorems concern the mean field limit of the ensemble Kalman filter; it is of interest to study finite particle approximations of the mean field limit, along the lines of the work in [44, 6] and analogous continuous time analyses in [17, 18, 19]. Analysis of mean field limits of interacting particle systems is an established field; interfacing the natural metrics employed for such analyses (Wasserstein) with those employed here (weighted TV) will be required.
@@
    We have made strong assumptions on $\Phi(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ and $h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ in this paper; developing proofs which relax these assumptions, is important. It will likely require use of Lyapunov functions and lead to bounds which hold in high probability with respect to the observed data.
@@
    Our error bounds are for a finite number of steps and, as is typical of finite time error estimates that employ a consistency plus stability implies convergence approach, lead to error constants which grow exponentially with the time horizon. Generalizing to long-time estimates, using assumptions on the long-term stability, or even ergodicity, of the true filter, would be of great interest. Related issues are widely studied for the true filtering distribution; see [49, 15, 62, 61, 10, 53] for example. In the context of the EnKF they are investigated in the paper [13] which identifies linear and Gaussian filtering problems in which it is possible to generalize the large particle asymptotic analyses of [44, 46]. There is also analysis of the EnKF for nonlinear non-Gaussian problems, such as the data assimilation problem for the Navier-Stokes equation, but this work concerns only accuracy of state estimation, not the entire filtering distribution [39].
@@
    We have studied a (widely used) version of the ensemble Kalman filter which employs a specific transport map to approximate the conditioning step in the filter. Other transport maps are also used in practice, such as that leading to the ensemble square root filter; see as outlined in [8, Section 2]. Analyzing these other methods would be of great interest.
@@
    Filtering can be used to solve inverse problems, as outlined in [8, Section 4]; in particular some of these methods rely on filtering over the infinite time horizon, the analysis of which will require new ideas such as in [21].
@@
    Continuous time versions of our analysis would be of interest, and their relationship to the Kushner–Stratonovich equation [11]; the paper [13], which studies mean field limits of the ensemble Kalman–Bucy filter, may be important in this context; furthermore study of the ensemble Kalman sampler [27, 28, 9] for inverse problems, would be of interest.
@@
    The papers [44] and [19] propose reweighting of ensemble Kalman methods to make them unbiased; further analysis of this idea, and the development of new methods that can carry this out in a derivative-free fashion, would be of interest.
@@
    Ensemble Kalman filters in practice employ very small numbers of ensemble members, and use covariance (spatial) localization when the unknown states are fields; developing analyses which account for low rank approximations and localization would be a valuable step for the field. Furthermore it would also be of interest to study the problem of quantization of measures [31] in the context of filtering; however this is a very hard problem even in simple low dimensional settings [7] and studying it for the evolution of the filtering distribution will require substantial new ideas.
@@
    Particle filters suffer from a curse of dimensionality on certain families of high-dimensional problems [4, 56]. Studying ensemble Kalman methods, to determine whether they ameliorate this issue in the near-Gaussian setting, would be of interest. There are a variety of different models that can be employed to characterize families of high dimensional filtering problems [52] and such structural assumptions will undoubtedly affect the results that might be obtained for ensemble Kalman methods.
@@ Acknowledgments
JAC was supported by the Advanced Grant Nonlocal-CPD (Nonlocal PDEs for Complex Particle Dynamics: Phase Transitions, Patterns and Synchronization) of the European Research Council Executive Agency (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 883363). JAC was also partially supported by the Engineering and Physical Sciences Research Council (EPSRC) under grants EP/T022132/1 and EP/V051121/1. FH is supported by start-up funds at the California Institute of Technology. FH was also supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via project 390685813 - GZ 2047/1 - HCM. The work of AMS is supported by a Department of Defense Vannevar Bush Faculty Fellowship. UV is partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 810367), and by the Agence Nationale de la Recherche under grant ANR-21-CE40-0006 (SINEQ).
@@ Appendix A Auxiliary Results
Suppose that $X$ is a random variable with values in ${}^{d}$ and finite second moment, and let $m:=\mathbf{E}[X]$. Then
@@
The following lemma is very similar to [2, Lemma 3.1]; the only difference is that the weight on the left-hand side is given by $1+\lvert u\rvert^{2}$ instead of $u^{2}$. We give the proof for completeness.
@@
Let $g(u)=1+\lvert u\rvert^{2}$ and assume that $\mu_{1}$, $\mu_{2}$ are probability measures over ${}^{d}$ satisfying $\mu_{1}[g^{2}]<\infty$ and $\mu_{2}[g^{2}]<\infty$. Then, if $\mu_{1}\ll\mu_{2}$,
@@
$\lVert h\rVert_{\infty}\leqslant L_{\alpha}\lVert h\rVert_{1},\qquad h(y)=c_{1}g(y;m_{1},\mathit{S}_{1})-c_{2}g(y;m_{2},\mathit{S}_{2}).$ (A.4)
@@
Throughout this proof, $C$ denotes a constant whose value is irrelevant in the context and may change from line to line. The statement is obvious if $c_{1}c_{2}=0$, so we assume from now on that $c_{2}>0$. We divide the proof into three steps.
Step 1: simplification
Since both sides of the inequality (A.4) are invariant under translation, it is sufficient to consider the case where $m_{2}=0$. Note also that
$c_{1}g(y;m_{1},\mathit{S}_{1})-c_{2}g(y;0,\mathit{S}_{2})$
$\qquad=\frac{c_{2}}{\sqrt{\det\mathit{S}_{2}}}\left(\frac{c_{1}}{c_{2}}g\left(\sqrt{\mathit{S}_{2}^{-1}}y;\sqrt{\mathit{S}_{2}^{-1}}m_{1},\sqrt{\mathit{S}_{2}^{-1}}\mathit{S}_{1}\sqrt{\mathit{S}_{2}^{-1}}\right)-g\left(\sqrt{\mathit{S}_{2}^{-1}}y;0,\mathit{I}_{K}\right)\right),$
and so we can also assume without loss of generality that $c_{2}=1$ and $\mathit{S}_{2}=\mathit{I}_{K}$. Indeed, assume that the inequality (A.4) is satisfied for this particular case with a constant $\widetilde{L}_{\alpha}$. Since
@@
Therefore, using the change of variable $y\mapsto\sqrt{\mathit{S}_{2}^{-1}}y$ together with (A.4) in the aforementioned particular case, we obtain that
@@
$\lVert c_{1}g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m_{1},\mathit{S}_{1})-c_{2}g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{S}_{2})\rVert_{\infty}$
@@
$\qquad=\frac{\lvert c_{2}\rvert}{\sqrt{\det\mathit{S}_{2}}}\left\lVert\frac{c_{1}}{c_{2}}g\left(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\sqrt{\mathit{S}_{2}^{-1}}m_{1},\sqrt{\mathit{S}_{2}^{-1}}\mathit{S}_{1}\sqrt{\mathit{S}_{2}^{-1}}\right)-g\left(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{I}_{K}\right)\right\rVert_{\infty}$
@@
$\qquad\leqslant\frac{\widetilde{L}_{\alpha^{2}}\lvert c_{2}\rvert}{\sqrt{\det\mathit{S}_{2}}}\left\lVert\frac{c_{1}}{c_{2}}g\left(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\sqrt{\mathit{S}_{2}^{-1}}m_{1},\sqrt{S_{2}^{-1}}S_{1}\sqrt{S_{2}^{-1}}\right)-g\left(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{I}_{K}\right)\right\rVert_{1}$
@@
$\qquad=\frac{\widetilde{L}_{\alpha^{2}}}{\sqrt{\det\mathit{S}_{2}}}\left\lVert c_{1}g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m_{1},\mathit{S}_{1})-c_{2}g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{S}_{2})\right\rVert_{1}\leqslant\alpha^{K/2}\widetilde{L}_{\alpha^{2}}\left\lVert c_{1}g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m_{1},\mathit{S}_{1})-c_{2}g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{S}_{2})\right\rVert_{1},$

### v6

@@ 4 Discussion and Future Directions
We have provided the first analysis of the error incurred by ensemble Kalman filters, as approximations of true filtering distribution, beyond the linear Gaussian setting. We have employed an appropriate weighted TV metric and obtained new stability estimates in this metric, in order to establish the approximation results. Our framing of the problem is motivated by the framing of the analysis of the particle filter contained in [MR3375889, Section 1], a systematization of the original proof of convergence of the particle filter appearing in [del1997nonlinear]. Although it introduces new methodology and theoretical results for nonlinear Kalman filters, our work leaves open numerous avenues for further analysis; we now highlight those that we identify as particularly important. These remaining open problems are substantial, but the framework we map out in this paper is an appropriate one in which to address them.
@@
    Our theorems concern the mean field limit of the ensemble Kalman filter; it is of interest to study finite particle approximations of the mean field limit, along the lines of the work in [le2009large, MR2860672] and analogous continuous time analyses in [ding2021ensemble, ding2021ensembleb, ding2020ensemble]. Analysis of mean field limits of interacting particle systems is an established field; interfacing the natural metrics employed for such analyses (Wasserstein) with those employed here (weighted TV) will be required.
@@
    We have made boundedness assumptions on $\Psi(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ and $h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ in this paper. Developing proofs which relax these assumptions, allowing consideration of small nonlinear perturbations of the Kalman filter setting, for example, will be very valuable.
@@
    Our error bounds are for a finite number of steps and, as is typical of finite time error estimates that employ a consistency plus stability implies convergence approach, lead to error constants which grow exponentially with the time horizon. The literature on analysis of numerical methods for non-autonomous dynamical systems demonstrates that going beyond finite time error estimates that grow in time is, in general, not possible [stuart1998dynamical]; in that context, generalizing to long-time estimates requires assumptions on the long-term stability, or even ergodicity, of the dynamical system. Such long-term stability issues are widely studied for the true filtering distribution – see [ocone1996asymptotic, del2001stability, van2009uniform, tong2012ergodicity, crisan2020stable, sanz2015long] for example; they are complicated by the fact that the nonlinear evolution equation for the filtering distribution is non-autonomous due to the observation signal. In the context of the EnKF such stability estimates are used in the paper [MR3784489] which identifies linear and Gaussian filtering problems in which it is possible to generalize the large particle asymptotic analyses of [le2009large, mandel2011convergence]. There is also analysis of the EnKF for nonlinear non-Gaussian problems, such as the data assimilation problem for the Navier-Stokes equation, but this work concerns only accuracy of state estimation, not the entire filtering distribution [kelly2014well, biswas2024unified].
@@
    We have studied a (widely used) version of the ensemble Kalman filter which employs a specific transport map to approximate the conditioning step in the filter. Other transport maps are also used in practice, such as that leading to the ensemble square root filter; as outlined in [2022arXiv220911371C, Section 2]. Analyzing these other methods would be of great interest.
@@
    Filtering can be used to solve inverse problems, as outlined in [2022arXiv220911371C, Section 4]; in particular some of these methods rely on filtering over the infinite time horizon, the analysis of which will require new ideas such as in [durmus2020elementary].
@@
    Continuous time versions of our analysis would be of interest, and their relationship to the Kushner–Stratonovich equation [cricsan1999interacting]; the paper [MR3784489], which studies mean field limits of the ensemble Kalman–Bucy filter, may be important in this context; furthermore study of the ensemble Kalman sampler [garbuno2020interacting, garbuno2020affine, carrillo2019wasserstein] for inverse problems, would be of interest.
@@
    The papers [le2009large] and [ding2020ensemble] propose reweighting of ensemble Kalman methods to make them unbiased; further analysis of this idea, and the development of new methods that can carry this out in a derivative-free fashion, would be of interest.
@@
    Ensemble Kalman filters in practice employ very small numbers of ensemble members, and use covariance (spatial) localization when the unknown states are fields; developing analyses which account for low rank approximations and localization would be a valuable step for the field. Furthermore it would also be of interest to study the problem of quantization of measures [graf2000foundations] in the context of filtering; however this is a very hard problem even in simple low dimensional settings [caglioti2016quantization] and studying it for the evolution of the filtering distribution will require substantial new ideas.
@@
    Particle filters suffer from a curse of dimensionality on certain families of high-dimensional problems [MR2459233, snyder2008obstacles]. Studying ensemble Kalman methods, to determine whether they ameliorate this issue in the near-Gaussian setting, would be of interest. There are a variety of different models that can be employed to characterize families of high dimensional filtering problems [2018arXiv181006191S] and such structural assumptions will undoubtedly affect the results that might be obtained for ensemble Kalman methods.
@@ Acknowledgments
We are grateful to the reviewers for very useful comments on a previous version of this paper. In particular, we thank the reviewers for pointing out the elementary inequality in Lemma 13, and for showing us how this inequality could be used to significantly improve the proofs of Lemmas 14, 16 and 32. Notably, the constructive proof of Lemma 14 was proposed by a reviewer. JAC was supported by the Advanced Grant Nonlocal-CPD (Nonlocal PDEs for Complex Particle Dynamics: Phase Transitions, Patterns and Synchronization) of the European Research Council Executive Agency (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 883363). JAC was also partially supported by the Engineering and Physical Sciences Research Council (EPSRC) under grants EP/T022132/1 and EP/V051121/1. FH is supported by start-up funds at the California Institute of Technology. FH was also supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via project 390685813 - GZ 2047/1 - HCM. The work of AMS is supported by a Department of Defense Vannevar Bush Faculty Fellowship, and by the SciAI Center, funded by the Office of Naval Research (ONR), under Grant Number N00014-23-1-2729. UV is partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 810367), and by the Agence Nationale de la Recherche under grants ANR-21-CE40-0006 (SINEQ) and ANR-23-CE40-0027 (IPSO).
[…]
@@ Appendix A Auxiliary Results
Suppose that $X$ is a random variable with values in d and finite second moment, and let $m:=\mathbf{E}[X]$. Then
@@
The following lemma is very similar to [MR2730330, Lemma 3.1]; the only difference is that the weight on the left-hand side is given by $1+\lvert u\rvert^{2}$ instead of $u^{2}$. We give the proof for completeness.
@@
Let $g(u)=1+\lvert u\rvert^{2}$ and assume that $\mu_{1}$, $\mu_{2}$ are probability measures over d satisfying $\mu_{1}[g^{2}]<\infty$ and $\mu_{2}[g^{2}]<\infty$. Then, if $\mu_{1}\ll\mu_{2}$,
@@
For all $d\in\mathbf{N}^{+}$ and $\alpha>0$, there is $C>0$ such that for all $x_{0},x_{1},m_{0},m_{1}\in^{d}$ and all symmetric positive definite matrices $S_{0},S_{1}\in^{d\times d}$ satisfying
$S_{0}\succcurlyeq\frac{1}{\alpha}I_{d},\qquad S_{1}\succcurlyeq\frac{1}{\alpha}I_{d},$
it holds that
$(2\pi)^{\frac{d}{2}}\bigl{\lvert}g_{0}(x_{0})-g_{1}(x_{1})\bigr{\rvert}$ $\leqslant\frac{\alpha^{\frac{1+d}{2}}}{\sqrt{\operatorname{e}}}\lvert x_{1}-x_{0}-m_{1}+m_{0}\rvert+\frac{\alpha^{1+\frac{d}{2}}}{\operatorname{e}}\lVert S_{1}-S_{0}\rVert$
$\qquad\qquad+\alpha^{d}\left\lvert\sqrt{\det S_{1}}-\sqrt{\det S_{0}}\right\rvert.$ (A.3)
Here $g_{0}$ and $g_{1}$ denote the densities of $\mathcal{N}(m_{0},S_{0})$ and $\mathcal{N}(m_{1},S_{1})$, respectively.
Proof.
For $s\in[0,1]$, let $x_{s}=(1-s)x_{0}+sx_{1},$ as well as $m_{s}=(1-s)m_{0}+sm_{1}$ and $S_{s}=(1-s)S_{0}+sS_{1}.$ Let us also introduce the non-normalized densities
$\widetilde{g}_{0}=\exp\left(-\frac{1}{2}\bigl{\lvert}x-m_{0}\bigr{\rvert}_{S_{0}}^{2}\right),\qquad\widetilde{g}_{1}=\exp\left(-\frac{1}{2}\bigl{\lvert}x-m_{1}\bigr{\rvert}_{S_{1}}^{2}\right),$
and let $\lambda(s)\coloneq\exp\left(-\frac{1}{2}\bigl{\lvert}x_{s}-m_{s}\bigr{\rvert}_{S_{s}}^{2}\right).$ It holds that
$\widetilde{g}_{1}(x_{1})-\widetilde{g}_{0}(x_{0})=\int_{0}^{1}\frac{\mathrm{d}\lambda}{\mathrm{d}s}(s)\,\mathrm{d}s$ $=-\int_{0}^{1}\Bigl{[}(x_{1}-x_{0}-m_{1}+m_{0})^{\mathsf{T}}S_{s}^{-1}(x_{s}-m_{s})$
$\qquad\qquad\quad-\,\frac{1}{2}(x_{s}-m_{s})^{\mathsf{T}}S_{s}^{-1}(S_{1}-S_{0})S_{s}^{-1}(x_{s}-m_{s})\Bigr{]}\lambda(s)\,\mathrm{d}s.$
Therefore, using that $\max_{z\in}\left\lvert z\operatorname{e}^{-\frac{z^{2}}{2}}\right\rvert=\frac{1}{\sqrt{\operatorname{e}}}$ and $\max_{z\in}\left\lvert z^{2}\operatorname{e}^{-\frac{z^{2}}{2}}\right\rvert=\frac{2}{\operatorname{e}}$, we deduce that
$\bigl{\lvert}\widetilde{g}_{1}(x_{1})-\widetilde{g}_{0}(x_{0})\bigr{\rvert}$ $\leqslant\int_{0}^{1}\Bigl{[}\lvert x_{1}-x_{0}-m_{1}+m_{0}\rvert_{S_{s}}\left\lvert x_{s}-m_{s}\right\rvert_{S_{s}}$
$\qquad\qquad+\frac{1}{2}\left\lVert S_{s}^{-{\frac{1}{2}}}(S_{1}-S_{0})S_{s}^{-{\frac{1}{2}}}\right\rVert\left\lvert x_{s}-m_{s}\right\rvert_{S_{s}}^{2}\Bigr{]}\lambda(s)\,\mathrm{d}s$
$\leqslant\sqrt{\frac{\alpha}{\operatorname{e}}}\lvert x_{1}-x_{0}-m_{1}+m_{0}\rvert+\frac{\alpha}{\operatorname{e}}\lVert S_{1}-S_{0}\rVert.$
Applying the triangle inequality
$\bigl{\lvert}g_{0}(x_{0})-g_{1}(x_{1})\bigr{\rvert}\leqslant\frac{\bigl{\lvert}\widetilde{g}_{0}(x_{0})-\widetilde{g}_{1}(x_{1})\bigr{\rvert}}{\sqrt{(2\pi)^{d}\det S_{0}}}+\left\lvert\frac{1}{\sqrt{(2\pi)^{d}\det S_{0}}}-\frac{1}{\sqrt{(2\pi)^{d}\det S_{1}}}\right\rvert\bigl{\lvert}\widetilde{g}_{1}(x_{1})\bigr{\rvert},$
we obtain (A.3), which concludes the proof. ∎
Lemma 14.
@@
$\lVert\mathfrak{h}\rVert_{\infty}\leqslant L_{\alpha}\lVert\mathfrak{h}\rVert_{1},\qquad\mathfrak{h}(y)=c_{1}g(y;m_{1},\mathit{S}_{1})-c_{2}g(y;m_{2},\mathit{S}_{2}).$ (A.5)
@@
For conciseness, let $g_{1}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})=g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m_{1},\mathit{S}_{1})$ and $g_{2}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})=g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m_{2},\mathit{S}_{2}).$
Step 1. Simplification
It is sufficient to prove the statement for $c_{1}=c_{2}=1$. Indeed, suppose that there is $\widetilde{L}_{\alpha}>0$ such that
$\forall(m_{1},\mathit{S}_{1})\in^{K}\times\mathcal{S}^{K}_{\alpha},\qquad\forall(m_{2},\mathit{S}_{2})\in^{K}\times\mathcal{S}^{K}_{\alpha},\qquad\lVert g_{1}-g_{2}\rVert_{\infty}\leqslant\widetilde{L}_{\alpha}\lVert g_{1}-g_{2}\rVert_{1}.$ (A.6)
Then, by the triangle inequality, it holds that
$\bigl{\lVert}c_{1}g_{1}-c_{2}g_{2}\bigr{\rVert}_{\infty}$ $\leqslant\lvert c_{1}\rvert\bigl{\lVert}g_{1}-g_{2}\bigr{\rVert}_{\infty}+\lvert c_{1}-c_{2}\rvert\bigl{\lVert}g_{2}\bigr{\rVert}_{\infty}$
$\leqslant\lvert c_{1}\rvert\widetilde{L}_{\alpha}\bigl{\lVert}g_{1}-g_{2}\bigr{\rVert}_{1}+\lvert c_{1}-c_{2}\rvert\bigl{\lVert}g_{2}\bigr{\rVert}_{\infty}$
$\leqslant\widetilde{L}_{\alpha}\bigl{\lVert}c_{1}g_{1}-c_{2}g_{2}\bigr{\rVert}_{1}+\lvert c_{2}-c_{1}\rvert\bigl{\lVert}g_{2}\bigr{\rVert}_{1}+\lvert c_{2}-c_{1}\rvert\bigl{\lVert}g_{2}\bigr{\rVert}_{\infty}.$
By Jensen’s inequality, it holds that
$\lvert c_{1}-c_{2}\rvert=\left\lvert\int_{{}^{d}}c_{1}g_{1}(x)-c_{2}g_{2}(x)\,\mathrm{d}x\right\rvert\leqslant\bigl{\lVert}c_{1}g_{1}-c_{2}g_{2}\bigr{\rVert}_{1},$
and so we deduce that
$\bigl{\lVert}c_{1}g_{1}-c_{2}g_{2}\bigr{\rVert}_{\infty}\leqslant\bigl{\lVert}c_{1}g_{1}-c_{2}g_{2}\bigr{\rVert}_{1}\left(\widetilde{L}_{\alpha}+1+\sqrt{\frac{\alpha^{K}}{(2\pi)^{K}}}\right).$
Furthermore, since both sides of the inequality (A.6) are invariant under translation, it is sufficient to consider the case where $m_{2}=0$, which we do from now on. Finally, note that
$g_{1}(y)-g_{2}(y)=\frac{1}{\sqrt{\det\mathit{S}_{2}}}\left(g\left(\sqrt{\mathit{S}_{2}^{-1}}y;\sqrt{\mathit{S}_{2}^{-1}}m_{1},\sqrt{\mathit{S}_{2}^{-1}}\mathit{S}_{1}\sqrt{\mathit{S}_{2}^{-1}}\right)-g\left(\sqrt{\mathit{S}_{2}^{-1}}y;0,\mathit{I}_{K}\right)\right),$
and so we can also assume without loss of generality that $\mathit{S}_{2}=\mathit{I}_{K}$. Indeed, assume that the inequality (A.6) is satisfied in this particular case with a constant $\widehat{L}_{\alpha}$. Since
@@
Therefore, using the change of variable $y\mapsto\sqrt{\mathit{S}_{2}^{-1}}y$ together with (A.6) in the particular case $\mathit{S}_{2}=\mathit{I}_{K}$, we obtain that
@@
$\lVert g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m_{1},\mathit{S}_{1})-g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{S}_{2})\rVert_{\infty}$
@@
$\qquad=\frac{1}{\sqrt{\det\mathit{S}_{2}}}\left\lVert g\left(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\sqrt{\mathit{S}_{2}^{-1}}m_{1},\sqrt{\mathit{S}_{2}^{-1}}\mathit{S}_{1}\sqrt{\mathit{S}_{2}^{-1}}\right)-g\left(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{I}_{K}\right)\right\rVert_{\infty}$
@@
$\qquad\leqslant\frac{\widehat{L}_{\alpha^{2}}}{\sqrt{\det\mathit{S}_{2}}}\left\lVert g\left(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\sqrt{\mathit{S}_{2}^{-1}}m_{1},\sqrt{S_{2}^{-1}}S_{1}\sqrt{S_{2}^{-1}}\right)-g\left(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{I}_{K}\right)\right\rVert_{1}$
@@
$\qquad=\frac{\widehat{L}_{\alpha^{2}}}{\sqrt{\det\mathit{S}_{2}}}\left\lVert g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m_{1},\mathit{S}_{1})-g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{S}_{2})\right\rVert_{1}\leqslant\alpha^{\frac{K}{2}}\widehat{L}_{\alpha^{2}}\left\lVert g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m_{1},\mathit{S}_{1})-g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{S}_{2})\right\rVert_{1},$